40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  TCS elog  Not logged in ELOG logo
Message ID: 79     Entry time: Tue Jul 27 08:31:10 2010
Author: Aidan 
Type: Electronics 
Category: SLED 
Subject: Long term SLED test - Day 1 

 The measurement from the on-board PD of the Superlum SLED seems to be falling. This effect started around 5PM last night which is right about the time we moved the position of the PD that the SLED is illuminating on the optical table (via optical fiber).

Curiously, the current set point and delivered current to the SLED are dropping as well. 

Attachment 1: SLED_superlum_long_term_test_0002A.pdf  7 kB  Uploaded Tue Jul 27 09:32:52 2010  | Hide | Hide all
SLED_superlum_long_term_test_0002A.pdf
ELOG V3.1.3-